TOF Ion Beam Test at CERN SPS, October 22-November 3, 2003

Runs
Types
Formats


Table 1: Runs

Run Events Type Day Start Stop RICH run Tracker run Note
200 24057 1 24/10/2003 17:44 17:57 518 162 Beam settings: A/Z=2. No EventID.
201 0 1 24/10/2003 18:03 18:43 519 163 EventID OK. Data file deleted!
202 111536 1 24/10/2003 19:21 19:54 520 164
203 55349 1 24/10/2003 20:48 21:30
165
204 17675 1 24/10/2003 21:46 22:27 521 166
205 85246 1 24/10/2003 22:41 23:18 522 167
206 73738 1 25/10/2003 23:36 00:11 523 168 Black stuff on C1 to avoid light.
207 32736 1 25/10/2003 02:43 03:00 525 169
208 21104 1 25/10/2003 03:54 04:06 525 170 RICH overwrites run 525.
209 54443 1 25/10/2003 04:54 05:23 525 171 RICH overwrites run 525.
210 68401 1 25/10/2003 05:54 06:28 526 172
211 62337 1 25/10/2003 06:52 07:28 527 173 Problems with the primary beam.
212 37574 1 25/10/2003 07:33 08:15 529 174
213 4552 1 25/10/2003 08:29 08:31
175
214 10000 pedestal 25/10/2003 08:33


Pedestals (10'000 ev.).
215 10000 pedestal 25/10/2003 08:41


Pedestals (10'000 ev.).
216 15732 2 25/10/2003 08:59 09:10 530 176 Problems with the primary beam.
217 0 2 25/10/2003 09:15 09:38 530 177 RICH overwrites run 530 (?).
218 10000 pedestal 25/10/2003 09:48


Pedestals (10'000 ev.).
219 45745 2 25/10/2003 10:04 10:35 531 179
220 48052 2 25/10/2003 11:00 11:36 532 179
221 52392 2 25/10/2003 11:52 12:27 533 180
222 10000 pedestal 25/10/2003 13:19


Pedestals (10'000 ev.).
223 16920 2 25/10/2003 14:07 14:19 534 181
224 4294 3 25/10/2003 14:27 14:30

C1DU, C1DD, C2DL, C2DR OK. Others no.
225 44003 3 25/10/2003 14:56 15:31 535 182 Second ADC is not working.
226 10000 pedestal 25/10/2003 17:25


Pedestals (10'000 ev.).
227 3597 3 25/10/2003 18:38 18:42 536 183 Exchanged the gates of the two ADC: nothing changes.
228 3381 3 25/10/2003 18:43 18:44 536 183
229 32012 3 25/10/2003 18:50 19:14 536 183
230 10000 pedestal 25/10/2003 19:15


Pedestals (10'000 ev.).
231 22834 3 25/10/2003 19:34 19:51 539 184
232 20031 4 25/10/2003 19:52 20:07 539 184 Clock (250 ns) OK.
233 29364 4 25/10/2003 20:46 21:09 540 185 C2AL has problems: bad ADC channel.
234 10000 pedestal 25/10/2003 21:32


Pedestals (10'000 ev.).
235 47184 4 25/10/2003 21:37 22:10 541 186
236 52361 4 25/10/2003 23:08 23:46 542 187
237 54116 4 26/10/2003 00:25 00:57 543 188
238 26115 4 26/10/2003 01:46 02:01 544 189
239 10000 pedestal 26/10/2003 02:16


Pedestals (10'000 ev.).
240 7858 5 26/10/2003 02:24 02:35 546 191
241 6498 5 26/10/2003 03:59 04:04 547 192
242 48760 5 26/10/2003 04:38 05:08 547 193 RICH overwrites run 547 (?).
243 53447 5 26/10/2003 05:34 06:04 548 194
244 46880 5 26/10/2003 06:28 06:57 549 195
245 10000 pedestal 26/10/2003 07:22


Pedestals (10'000 ev.).
246 60497 6 26/10/2003 07:26 07:52 550 196
247 10000 pedestal 26/10/2003 08:00


Pedestals (10'000 ev.).
248 78041 6 26/10/2003 09:39 10:11 551 197
249 10000 pedestal 26/10/2003 10:33


Pedestals (10'000 ev.).
250 71601 6 26/10/2003 10:38 11:09 552 198
251 10000 pedestal 26/10/2003 11:24


Pedestals (10'000 ev.).
252 12421 6 26/10/2003 11:34 11:39 553 199
253 3918 6 26/10/2003 11:44 11:46 553 199 New ADC for dynodes. Not working.
254 3122 6 26/10/2003 11:48 11:49 553 199 Another ADC. Not working
255 3996 7 26/10/2003 11:52 11:55 553 199 Second ADC not working.
256 151 7 26/10/2003 11:56 11:58 553 199 Still not working!
257 22174 8 26/10/2003 11:59 12:07 553 199
258 106376 8 26/10/2003 12:38 13:21 554 200
259 96103 9 26/10/2003 15:42 16:20 555 202
260 10000 pedestal 26/10/2003 16:54


Pedestals (10'000 ev.).
261 67628 9 26/10/2003 16:59 17:26 557 206
262 10000 pedestal 26/10/2003 18:24


Pedestals (10'000 ev.).
263 75441 9 26/10/2003 20:49 21:22 561 207
264 10000 pedestal 26/10/2003 21:54


Pedestals (10'000 ev.).
265 19686 9 26/10/2003 22:08 22:19
208
266 74879 9 26/10/2003 22:20 22:54 562 208
267 10000 pedestal 26/10/2003 23:10


Pedestals (10'000 ev.).
268 11361 9 26/10/2003 23:20 23:25 563 209 Beam power failure.
269 75639 9 26/10/2003 23:34 00:05 564 210
270 85289 9 27/10/2003 00:24 01:03 565 211
271 36151 9 27/10/2003 01:30 01:46 566 212
272 62794 9 27/10/2003 02:14 02:39 567 213
273 10000 pedestal 27/10/2003 02:42


Pedestals (10'000 ev.).
274 13147 9 27/10/2003 03:03 03:10 568 214
275 35220 9 27/10/2003 03:34 03:50 569 215
276 6562 9 27/10/2003 04:19 04:27 570 216 Beam problems.
277 56544 9 27/10/2003 05:40 06:14 571 217
278 10000 pedestal 27/10/2003 06:22


Pedestals (10'000 ev.).
279 27962 9 27/10/2003 06:35 06:54 572 218 Interrupted and started again after 574 ev.
280 10354 9 27/10/2003 07:19 07:26 573 219
281 17382 9 27/10/2003 07:47 07:56 574 220
282 841 10 27/10/2003 09:40 09:45


283 10000 pedestal 27/10/2003 09:48


Pedestals (10'000 ev.).
284 41906 10 27/10/2003 10:00 10:37 575 221
285 10000 pedestal 27/10/2003 10:45


Pedestals (10'000 ev.).
286 5490 11 27/10/2003 12:00 12:18 576 222 Beam problems.
287 10000 pedestal 27/10/2003 12:25


Pedestals (10'000 ev.).
288 4438 11 27/10/2003 13:22 13:30 577 223 Beam problems.
289 10000 11 27/10/2003 13:31
578 224 Only 10'000 ev.
290 55387 11 27/10/2003 14:00 14:38

RICH & tracker not running.
291 31543 11 27/10/2003 14:39 15:08 579 225
292 10000 pedestal 27/10/2003 15:23


Pedestals (10'000 ev.).
293 35043 11 27/10/2003 15:37 16:03 580 226
294 42344 11 27/10/2003 16:22 16:51 581 227
295 10000 pedestal 27/10/2003 16:52


Pedestals (10'000 ev.).
296 1000 pedestal 27/10/2003 18:52


Pedestals (10'000 ev.).
297 36941 12 27/10/2003 18:57 19:26 583 229
298 36624 12 27/10/2003 19:46 20:46 584 230
299 16118 12 27/10/2003 20:49 21:12 585 231
300 33134 13 28/10/2003 02:05 02:34 590 240 Gate 1° ADC not connected! Bad trigger condition!
301 10000 pedestal 28/10/2003 02:50


Pedestals (10'000 ev.). ???
302 18389 13 28/10/2003 02:53 03:16 591 241
303 9173 13 28/10/2003 05:18 05:48 592 243
304 28941 13 28/10/2003 07:15 07:45 594 245
305 26332 13 28/10/2003 08:05 08:35 595 246 Beam problems.
306 76254 13 28/10/2003 10:30 11:17 599 247
307 82499 14 28/10/2003 13:12 14:17 600 248 >>> A/Z = 2.25 (heavy ions enriched beam).
308 62371 14 28/10/2003 14:40 15:40 601 249 Changed clock into the scaler: 185 microsec.
309 10000 pedestal 28/10/2003 15:54


Pedestals (10'000 ev.).
310 16054 15 28/10/2003 16:36 16:49 602 252
311 149116 16 28/10/2003 17:04 18:05 603 254 Luminosity increased during this run.
312 10000 pedestal 28/10/2003 18:22



313 70863 17 28/10/2003 18:30 19:07 604/6 255
314 73953 17 28/10/2003 19:08 19:32
256
315 149165 pedestal 28/10/2003 20:41 22:00

No veto in trigger! Bad run.
316 155 18 28/10/2003 22:13 22:18

idem
317 20635 18 28/10/2003 22:37 22:40

idem
318 10000 pedestal 28/10/2003 22:44


Pedestals (10'000 ev.).
319 14414 18 28/10/2003 22:58 23:06 608 260
320 38870 18 28/10/2003 23:58 00:06

RICH stopped DAQ: no EventID.
321 206793 18 29/10/2003 00:30 01:07

idem
322 217583 18 29/10/2003 01:10 01:50

idem
323 2397118 18 29/10/2003 02:01 09:53

idem
324 10000 pedestal 29/10/2003 09:54


Pedestals (10'000 ev.).
325 10000 pedestal 30/10/2003 09:30


Pedestals (10'000 ev.).
326 10000 pedestal 30/10/2003 10:59


Pedestals (10'000 ev.).
327 10000 pedestal 30/10/2003 11:10


Pedestals (10'000 ev.).
328 31374 19 30/10/2003 11:15


In beam.
329 87528 19 30/10/2003 11:32 11:41 612 267 In beam.
330 154925 19 30/10/2003 12:38 13:35 613 269 In beam and A/Z = 2.25 (?).
331 55981 19 30/10/2003 13:41 13:53 614 270
332 31754 19 30/10/2003 13:53 15:27 615

333 10000 pedestal 30/10/2003 15:33


Pedestals (10'000 ev.).
334 87210 20 30/10/2003 15:44 16:11 615 274
335 100929 20 30/10/2003 16:13 16:42 615 274
336 10000 pedestal 30/10/2003 16:44


Pedestals (10'000 ev.).
337 100957 20 30/10/2003 16:46 17:13 616 276/7 Beam off.
338 10000 pedestal 30/10/2003 17:15


Pedestals (10'000 ev.).
339 38499 20 30/10/2003 21:20 21:41
284 Protons 5 GeV.
340 10000 pedestal 30/10/2003 21:43


Pedestals (10'000 ev.).
341 68257 20 30/10/2003 21:45 22:25 621 286 Protons 9 GeV.
342 10000 pedestal 30/10/2003 22:28


Pedestals (10'000 ev.).
343 9568 20 30/10/2003 22:32 22:55 622 287 Protons 7 GeV.
344 10000 pedestal 30/10/2003 22:57


Pedestals ?
345 10000 pedestal 30/10/2003 22:53


Pedestals (10'000 ev.).
346 14220 20 30/10/2003 23:01 23:15 622 287 Trigger rate low: tests during this run.
347 91574 21 31/10/2003 00:48 01:14 624 288/9 Protons 15 GeV.
348 10000 pedestal 31/10/2003 01:14


Pedestals (10'000 ev.).
349 38253 21 31/10/2003 01:28 01:53 625
Beam tuning near the end of run.
350 10000 pedestal 31/10/2003 02:31


Pedestals (10'000 ev.).
351 99860 21 31/10/2003 02:33 03:28 627 291 Protons 9 GeV.
352 10000 pedestal 31/10/2003 03:29


Pedestals (10'000 ev.).
353 61903 21 31/10/2003 03:35 04:58 628 292 Protons 7 GeV.
354 10000 pedestal 31/10/2003 04:59


Pedestals (10'000 ev.).
355 96515 22 31/10/2003 05:30 06:12

Beam off.
356 10000 pedestal 31/10/2003 06:12


Pedestals (10'000 ev.).
357 113991 23 31/10/2003 06:15



358 10000 pedestal 31/10/2003 06:56


Pedestals (10'000 ev.).
359 90304 24 31/10/2003 06:57 07:25


360 93341 24 31/10/2003 07:28 07:57


361 10000 pedestal 31/10/2003 07:58


Pedestals (10'000 ev.).
362 93587 24 31/10/2003 08:03 08:35 629 294
363 10000 pedestal 31/10/2003 10:32


Pedestals (10'000 ev.).
364 90391 25 31/10/2003 10:32 11:09 630 295
365 10000 pedestal 31/10/2003 11:10


Pedestals (10'000 ev.).
366 32890 25 31/10/2003 11:12 11:24


367 10000 pedestal 31/10/2003 11:43


Pedestals (10'000 ev.).
368 74246 26 31/10/2003 11:44 12:07

Beam tuning on course.
369 356147 26 31/10/2003 12:45 14:07 631/2

370 18861 26 31/10/2003 14:08


Pedestals (10'000 ev.).
371 185293 26 31/10/2003 15:13 15:55 636/7

372 10000 pedestal 31/10/2003 15:56


Pedestals (10'000 ev.).
373 116378 27 31/10/2003 16:00 16:30


374 138855 27 31/10/2003 16:55 17:36 638

375 10000 pedestal 31/10/2003 17:37


Pedestals (10'000 ev.).
376 105888 27 31/10/2003 17:40 17:53 639

377 10000 pedestal 31/10/2003 17:54


Pedestals (10'000 ev.).
378 75288 27 31/10/2003 17:56 18:26 640

379 10000 pedestal 31/10/2003 18:25


Pedestals (10'000 ev.).
380 10000 pedestal 01/11/2003 09:27


Pedestals (10'000 ev.).
381 91616 28 01/11/2003 09:28 10:08


382 10000 pedestal 01/11/2003 10:11


Pedestals (10'000 ev.).
383 3201 29 01/11/2003 10:21 10:51

Changed configuration on course.
384 13058 30 01/11/2003 10:51 11:17

~100 particles/spill. No C3AL!
385 8721 30 01/11/2003 11:18 11:19

C3AL OK.
386 193405 31 01/11/2003 11:22 11:56


387 10000 pedestal 01/11/2003 11:56


Pedestals (10'000 ev.).
388 10000 pedestal 01/11/2003 12:08


Pedestals (10'000 ev.).
389 107891 32 01/11/2003 12:09 12:28


390 496101 32 01/11/2003 12:28 14:03


391 10000 pedestal 01/11/2003 14:03


Pedestals (10'000 ev.).
392 282598 33 01/11/2003 14:11 15:04


393 10000 pedestal 01/11/2003 15:05


Pedestals (10'000 ev.).
394 369938 33 01/11/2003 15:08 16:21


395 10000 pedestal 01/11/2003 16:22


Pedestals (10'000 ev.).
396 282629 33 01/11/2003 16:34 17:31


397 191893 33 01/11/2003 17:31 18:10


398 10000 pedestal 01/11/2003 18:10


Pedestals (10'000 ev.).

Table 2: Types

Type First run Format Explanation
1 200 1 Nominal voltages. C3 and C4 without dynodes. A/Z = 2 beam.
2 215 1 Turned off half of PMTs: 5554, 5552, 7083, 5551, 5893, 5915, 5974, 6012
3 224 2 Connected C1DU, C1DD, C2DL, C2DR to the anodes ADC.
4 232 3 Last scaler channel always zero: no clock. Added dummy channel to recover 250 ns clock.
5 240 3 Hour change: from 02:59:59 to 02:00:00.
6 245 3 All PMTs on.
7 255 4 Tested many dynodes configurations and added ADC modules.
8 257 5 C3DL, C3DR, C4DL, C4DR still not working.
9 259 5 Tested many dynodes configurations and used different ADC modules.
10 282 5 Turned off half of PMTs: 5555, 5553, 7077, 5542, 5869, 5970, 5886, 5975
11 286 5 Tested differentctions and addedof dynodes. Still not working.
12 296 6 Installed other ADC modules.
13 300 6 Changed long cable for C1DU. Still very low pulse.
14 307 6 A/Z = 2.25 (heavy ions enriched beam).
15 310 6 Changed clock from 250 ns to 185 us. Moved dynodes of C3 and C4 on the highest fraction of the passive divider.
16 311 7 Luminosity increase during run.
17 312 7 Adjusted zero of first linear fan-out. C1AU out from trigger. Both PMTs on C1up turned on (very low pulse).
18 315 7 Second fan-out dos not work. Plugged 1 dynode/counter in first fan-out.
19 325 7 Changed second linear fan-out. Recovered C3 and C4 dynodes.
20 333 7 All PMTs on.
21 347 7 Rate very low since run 343 with protons beam. Thresholds from 30 mV to 16 mV. C4AL out from the trigger (as C1AU).
22 355 7 RICH and tracker stopped DAQ. A/Z=2 beam.
23 357 7 C4AL again in trigger.
24 358 7 Thresholds from 16 mV to 30 mV.
25 363 7 Protons beam.
26 367 7 A/Z = 2.35 Indium beam.
27 371 7 The tracker stops running.
28 380 7 The RICH stops running. All PMTs on with nominal settings.
29 383 7 A/Z = 2 beam.
30 384 7 Rate very low. Thresholds at 16 mV.
31 386 7 Rate OK. Thresholds at 30 mV.
32 388 7 Changed HV for C2, C3 and C4. Their gain is 60% as before.
33 392 7 Thresholds at 16 mV. 60% gain.

Table 3: Formats

No. FORTRAN77 format Variables
1 8z4.4,4x,8z4.4,6x,1z6.6,11z4.4,8z4.4 Anodes, dynodes, EventID, scalers,times
2 8z4.4,8z4.4,6x,1z6.6,11z4.4,8z4.4 Anodes, dynodes, EventID, scalers,times
3 8z4.4,8z4.4,6x,1z6.6,11z4.4,4x,8z4.4 Anodes, dynodes, EventID, scalers,times
4 8z4.4,8z4.4,30x,1z6.6,11z4.4,4x,8z4.4 Anodes, dynodes, EventID, scalers,times
5 8z4.4,8z4.4,6x,1z6.6,11z4.4,4x,8z4.4 Anodes, dynodes, EventID, scalers,times
6 2z4.4,4x,6z4.4,4x,2z4.4,36x,3z4.4,2x,1z6.6,36x,3z4.4,11z4.4,4x,8z4.4 Anodes, C1& C2 dynodes, C3DL, EventID, C3DR, C4 dynodes, scalers, times.
7 2z4.4,4x,6z4.4,3z4.4,40x,2z4.4,36x,3z4.4,2x,1z6.6,11z4.4,4x,8z4.4 Anodes, C3 & C4 dynodes, C1 & C2 dynodes, EventID, scalers, times.